OpenStudy (anonymous):

find the vertex,focus and directrix of the parabola x^2+8y=0

OpenStudy (anonymous):

x^2+8y=0 ((x^2)- 0)+8(y-0)=0 so, vertex is (0,0) directrix is x=0

OpenStudy (anonymous):

directrix is actually y=2